How many milliliters of 3.00 M HCl(aq) are required to react with 6.35 g of an ore containing 28.0% Zn(s) by mass?
Answer #1

Zn (s) + 2 HCl (aq) -------> ZnCl2 (aq) + H2 (g)

mass of Zn = (6.35 *28 /100) = 1.778 g of Zn

number of moles of Zn = 1.778/ 65.39 = 0.02719 moles

1 mole of Zn requires 2 moles of HCl

number of moles of HCl = 0.02719*2 = 0.05438 moles

volume of HCl = (0.05438 / 3) L

                          = 18.13 mL of 3 M HCl is required